Carpenter Andersen owns a lumber company and to build a house. The whole house can be built in 31 days with five carpenters. Ten days after construction began, hired Andersen in 10 more carpenters to be previously finished. How many more days are needed then to finalization of the house?
Answer by stanbon(75887)

Carpenter Andersen owns a lumber company and wants to build a house. The whole house can be built in 31 days with five carpenters. Ten days after construction began, Andersen hired in 10 more carpenters. How many more days are needed then to finish the house?
----
# of man-days to complete the job = 31*5 = 155
After 10 days 5*10 = 50 man-days of the job are completed by the 5.
Ans: With 15 men working it will take 105/15 = 7 more days to complete the job
-------------------------
Cheers,
Stan H.
=========================